OpenStudy (anonymous):

Through (-2, 1) and parallel to y = -3x + 1?

OpenStudy (texaschic101):

a parallel line will have the same slope as the equation given. Can you tell me the slope of the equation given ?

OpenStudy (texaschic101):

in y = mx + b form, the slope is in the m position

rishavraj (rishavraj):

x+y+5=0

OpenStudy (texaschic101):

the slope of the equation y = -3x + 1...is -3 (because it is in the m position) now take our points (-2,1)..x = -2 and y = 1 slope(m) = -3 now sub into y = mx + b form y = mx + b 1 = -3(-2) + b 1 = 6 + b 1 - 6 = b -5 = b so your parallel equation is : y = -3x - 5